One side of a triangle is 2 cm shorter than the base. The other side is 3 cm longer than the base. What lengths of the base will allow the perimeter to be greater than 19 cm? Show your work and explain your answer.

Answer by stanbon(75887)

One side of a triangle is 2 cm shorter than the base. The other side is 3 cm longer than the base. What lengths of the base will allow the perimeter to be greater than 19 cm? Show your work and explain your answer.
-------------------
Draw the triangle.
Let the base be "x":
Let a shorter side be "x-2"
Let the 3rd side be "x+3".
------------------------------
Perimeter = 19 = x + x-2+x+3
19 = 3x+1
3x = 18
x = 6 cm (base)
x-2 = 4 cm (shorter side)
x+3 = 9 cm (longer side)
=============================
Cheers,
Stan H.